OpenStudy (anonymous):

Determine the per annum interest rate required for an investment with continuous compound interest to triple in 8 years. Express your answer as a percent. Enter an exact answer or round to three decimal places, for example 10.227%. The answer I got was ln3/8 but it says its wrong idk y?

OpenStudy (anonymous):

\[3=e^{8t}\] \[t=\frac{\ln(3)}{8}\] what you said

OpenStudy (anonymous):

maybe it does not like your rounding

OpenStudy (anonymous):

? ok cause that is wat I got

OpenStudy (anonymous):

or maybe it wants you to write \[\frac{\ln(3)}{8}\] and you entered \[\frac{\ln 3}{8}\] who knows?

OpenStudy (anonymous):

i tried both lol thanks anyway